OpenStudy (31356):

Karin wants to use the distributive property to find the value of 19 • 42 + 19 • 58 mentally. Which expression can she use? a)19*(42+58) b)(19+19)*(42+58) c)19*(42*58) d)19*19*(42+58) FREE MEDAL REWARDED!!

OpenStudy (anonymous):

d for sure

OpenStudy (31356):

Thx!

OpenStudy (poopsiedoodle):

It's really not D though. Just a second, let me work it out.

OpenStudy (31356):

I know I got the answer wrong

OpenStudy (31356):

can you help me with another problem?

OpenStudy (anonymous):

i am sorry i did it on my calculator and that is the only one that got the same answer as the actual problem

OpenStudy (31356):

That's okay

OpenStudy (anonymous):

well what is the other problem

OpenStudy (31356):

8*4=(triangle*7)-(triangle*3) What is the value of the triangle?

OpenStudy (poopsiedoodle):

The first one was A, since you're multiplying 19 by 100.

OpenStudy (poopsiedoodle):

Could you explain your second problem a tad bit more?

OpenStudy (31356):

Okay thx

OpenStudy (31356):

Just find the value of the triangle.

OpenStudy (31356):

Any ideas?

OpenStudy (poopsiedoodle):

Triangle = 8. 8*4 = 32. 32 = 8*7 (or 56) - 8*3 (which is 24). So, 32 = 4triangle. Divide both sides by 4, and you get 8 = triangle.

OpenStudy (31356):

Thx!

OpenStudy (31356):

Can you help me with this too?

OpenStudy (31356):

You and three friends own a paint shop. The shop's profit was $536 in the first month. You always divide the profits equally. In two months, each of you had made $224. How much profit did the shop make in the second month? a)$88 b)$360 c)$312 d)$1920

OpenStudy (anonymous):

hey @poopsiedoodle what is with the a by your name are u an administrator

OpenStudy (31356):

Any ideas?

OpenStudy (anonymous):

and @poopsiedoodle why did u pick poopsiedoodle

OpenStudy (poopsiedoodle):

So, if the profit the first month was $536, then you would each get $134. So, subtract that from the total of $224, and you get $90. Multiply that by 4, and you get $360. So, b.

OpenStudy (31356):

Thx!

OpenStudy (anonymous):

hey @poopsiedoodle what is with the a by your name are u an administrator

OpenStudy (31356):

Can you help me with this too?

OpenStudy (poopsiedoodle):

froggy, check your pms.

OpenStudy (31356):

What does n equal?

OpenStudy (31356):

Any ideas?

OpenStudy (poopsiedoodle):

n = 3. -4(3-6) gives you 12. 3-6 = -3, and when you multiply -4 by -3, you get positive 12 instead of -12.

OpenStudy (31356):

Okay thx

OpenStudy (anonymous):

is you are an ambassador then shouldn't u enforce the rule to lead geusts to the answer not give it to them

OpenStudy (poopsiedoodle):

yw :3

OpenStudy (31356):

Thx some much today!

OpenStudy (poopsiedoodle):

Froggy, I'm giving the answer and explaining, which, as far as I know, is ok with the mods.